OpenStudy (anonymous):

A temperature of 0 degrees C equals 32 degrees F and 25C equals 77F use this equation to find the fahrenheit temperature for 40C

OpenStudy (anonymous):

F=9C/5+32 I think

OpenStudy (anonymous):

To answer this, they're assuming you'll assume the relationship between the temperatures is linear. You want Fahrenheit to be the 'y' and Celsius to be the 'x'. You know the following: when x = 0, y = 32 ---> (x_1, y_1) when x = 25, y = 77 ---> (x_2, y_2) You may now use the two-point formula for a straight line: \[y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1) \rightarrow y-32=\frac{77-32}{25-0}(x-0)\] that is,\[y-32=\frac{45}{25}x \iff y=\frac{9}{5}x+32\]

OpenStudy (anonymous):

When the temperature is 40 degrees C, x=40, and so the temp. in F is,\[y=\frac{9}{5}40+32=104 ^oF\]

OpenStudy (anonymous):

Yeah, Rouault, you were right ;)

OpenStudy (anonymous):

:) I googled it before you answered to see